Status Report

NASA Space Station On-Orbit Status 1 February 2006

By SpaceRef Editor
February 1, 2006
Filed under , , ,
NASA Space Station On-Orbit Status 1 February 2006

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2006) to enhance access to related status reports and NASA activities. 

All ISS systems continue to function nominally, except those noted previously or below.

Onboard sleep cycle continues to be shifted 7 hours to the right (wake up 8:00am EST, sleep at 11:30pm), to prepare the crew for next Friday’s EVA-15 spacewalk (hatch opening: ~5:20pm EST).

FE Valery Tokarev’s regular morning inspection today included the routine inspection of DC-1 circuit breakers and fuses.   [The monthly checkup in the “Pirs” Docking Compartment looks at AZS circuit breakers on the BVP Amp Switch Panel (they should all be On) and the LEDs (light-emitting diodes) of 14 fuses in Fuse Panels BPP-30 & BPP-36.]

After their successful suited dry run exercise, the crew today continued preparations for the Orlan EVA-15 on 2/3.   [Yesterday’s dry run included Orlan systems checks (oxygen, power, ventilation, and cooling), EVA gear donning, Orlan communications check & medical parameter test, Orlan suit donning, Orlan fit checks, pressure & leak testing, suit doffing and post dry-run activities.  The dry run also re-acquainted the crew with the EVA tools to be used, the feel and functions of the Orlan, and operation of onboard systems in preparation for the EVA.  During the suit fit check, FE Tokarev reported extra pressure in the shoulder area of the suit and that his suit arms were too long.  CDR McArthur found that his suit arms were too short.  The crew depressurized and doffed their suits, made additional suit sizing changes and repressurized the suits.  The subsequent fit checks were successful, but the FE indicated that he might make minor adjustments in suit sizing prior to the EVA.]


Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

Tagging up with ground specialists, the crew reviewed specific EVA procedures such as steps for removing/retrieving the Russian “Biorisk” container #2 from the DC1 shell and procedures for effecting an emergency ingress in the Service Module Transfer Compartment (SM PkhO) in case of a leak in the DC1 EV1 hatch.  Afterwards, McArthur and Tokarev also reviewed U.S. EVA procedures and at ~1:45pm EST conducted a 30-min. spacewalk telecon with MCC-Houston via S-band.   [Originally considered only for Russian tasks, the ~6-hour spacewalk has become a bi-lateral EVA due to the late addition of the planned work on the MT (Mobile Transporter), consisting of installation of a safing bolt on its zenith IUA (Integrated Umbilical Assembly) after an uncommanded actuation of the nadir IUA “guillotine” resulted in severing the associated TUS (Trailing Umbilical System) cable; recent (1/13) CBM (Common Berthing Mechanism) surveys also discovered a missing launch restraint bolt on the zenith IUA which the crew will attempt to drive. The other (Russian) spacewalk objectives are: (a) Retrograde jettisoning of “RadioSkaf” (from “Radio” plus “Skafandr”), a worn-out Orlan suit containing ham radio equipment for young amateurs and transmission of a congratulatory message to the Moscow Bauman Institute in multiple languages, (b) relocation of the Russian/US Strela 2 crane adapter from the EFGF (Electrical Flight Grapple Fixture) on the FGB module to the FRGF (Flight Releasable GF) on the PMA-3 (Pressurized Mating Adapter #3), deferred from EVA-14 and required for Mission 12A.1 SMDP (SM Debris Panel) installation, (c) “Biorisk-2” retrieval, and (d) photo documentation of SMMK (micrometeoroid monitoring system) sensors, Soyuz TMA-7 Descent Module MLI (multi-layer insulation) and panorama surveys of the Russian segment (RS) exterior.]

The FE is performing closeout operations on the EVA hardware to be taken out.   [These include: preparation of a carrier bundle containing bungee tool caddy, tethers, plus dry towels for the “Biorisk” retrieval, configuring the KPU tool carrier and attaching it near the EV1 hatch, final set-up of RadioSkaf, and restoration of the Orlan systems after the suited dry run with fresh LP-9 lithium hydroxide (LiOH) canisters, new BK-3 oxygen repressurization tanks, refilled U.S. DIBS (disposable in-suit drink bags) as required, etc.]

Later today, the CDR will initiate charging of the DCS-760 digital camera to be used in the EVA and configure the camera.  After that, he will terminate the maintenance cycle of the EMU batteries in the Airlock and then transfer SODF (Systems Operations Data File) books and the CCPK (Crew Contamination Protection Kit) from the USOS to the SM.   [SODF items include hardcopies of the SODF Warning book, POC book and Medical Checklist, as well as portions of the CD library.] 

Late tonight, before sleep time, McArthur also will close the U.S. Airlock hatch preparatory to the EVA.

Valery Tokarev performed the daily routine maintenance of the SM’s environment control & life support system (SOZh), including the toilet system (ASU).

The FE is also scheduled to deactivate the SM’s IK0501 gas analyzer (GA) of the SOGS Pressure Control & Atmospheric Monitoring System and exchange its carbon dioxide (CO2) filter assembly (BF) with a new unit from FGB stowage (replaced last: 12/22/05).  GA will be reactivated and the spent BF stowed for disposal.  In addition, the FE is to clear the volume behind SM panel 449 of equipment blocking air flow and to take flash photography with the Nikon D1X for subsequent downlink.   [IK0501 is an automated system for measuring CO2, O2, and H2O in the air as well as the flow rate of the gas being analyzed.]

At the same time, Bill takes the weekly cabin air data with the GANK-4M Real-Time Harmful Contaminant Gas Analyzer system of the SOGS, which tests particularly for NH3 (ammonia) and HCl (hydrogen chloride).

Later, McArthur will perform the daily atmospheric status checks for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen sensor) and CDMK (CO2 Monitoring Kit).

Tokarev is to conduct the regular periodic download of data & imagery collected of the BIO-5 Rasteniya-2 (“Plants-2”) experiment to the computer for subsequent downlink to the ground.  Working off his discretionary task list, Valery will also conduct the daily status check on the greenhouse payload.   [Rasteniya researches growth and development of plants under spaceflight conditions in the Lada-8 greenhouse.  The regular maintenance of the experiment (each Monday, Wednesday, Friday and Sunday) involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, topping off the water tank if ~20-25% of the total amount (4 liters) remains, and photo/video recording.  Once weekly, data from the Lada greenhouse control unit are recorded on floppy disk for weekly downlink via REGUL-Packet or the new BSR-TM at a suitable occasion

At ~8:00pm EST, the CDR is scheduled for the weekly audit/inventory of the available CWCs (collapsible water containers) and their contents, to keep track of onboard water supplies.   [Updated “cue cards” based on Bill’s water calldowns are sent up every other week.]

The Science Officer is continuing his support of the PromISS-4 (Protein Crystal Growth Monitoring by Digital Holographic Microscope #4) experiment today by replacing tape #1 on the MSG (Microgravity Science Glovebox) video system with tape #2.   [Tape exchanges vary due to record times throughout the 15-day running period of the experiment.]

The crew has time reserved for their regular 2.5-hr. physical exercise program on the CEVIS cycle ergometer, TVIS treadmill and RED resistive exerciser.  Valery’s workout continues his prescribed four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 3 of a new set).

Afterwards, Bill transfers the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

McArthur’s schedule today also includes the standard weekly maintenance on the TVIS in the SM, primarily checking the condition of the SPDs (Subject Positioning Devices) and recording time & date values.

Later today, the FE will check the operation of the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at 20 degC, including a temperature check on its ART (automatic temperature recorder).   [This daily monitoring/temp checking, carried on the Russian voluntary “time available” task list, will continue until 4/30.]

Today’s CEO (Crew Earth Observation) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Internal waves, West Azores, Atlantic (clear weather was predicted following passage of a frontal system. Looking to the right of track [SW of the islands] for the sunglint point and internal wave patterns.  Referencing of internal wave images back to a geographic feature via overlapping frames is desired to aid in geolocation), Internal waves, North Patagonian Shelf (interesting internal wave patterns may be visible in embayments along the South American coastline.  Looking to the right of track for the sunglint point), and Calcutta, India (Night Time Practice Target.  The crew had a nadir pass over the center of this megacity).

Over 177,000 of CEO (Crew Earth Observation) images have been taken in the first five years of the ISS.

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 12 crew visit:

Expedition 12 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 7:27am EST [= epoch]):

  • Mean altitude — 346.3 km
  • Apogee height — 352.5 km
  • Perigee height — 340.1 km
  • Period — 91.46 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0009254
  • Solar Beta Angle — -36.3 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.74
  • Mean altitude loss in last 24 hours — 55 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 41178

Events Ahead (all dates Eastern; tentative):

  • 02/03/06 — Russian EVA-15
  • 02/11/06 — ISS Reboost Test (in MMOD avoidance mode)
  • 03/03/06 — Progress M-54/19P undocking & reentry
  • 03/30/06 — Soyuz TMA-8/12S launch (Exp. 13 + Marcos Pontes/Brazil)
  • 04/01/06 — Soyuz TMA-8/12S docking (DC1)
  • 04/09/06 — Soyuz TMA-7/11S undocking & reentry
  • 04/16/06 — Soyuz TMA-8/12S relocation (DC1 to FGB nadir port)
  • 04/24/06 — Progress M-56/21P launch
  • 04/26/06 — Progress M-56/21P docking
  • 06/19/06 — Progress M-55/20P undocking & reentry
  • 06/28/06 — Progress M-57/22P launch
  • 06/30/06 — Progress M-57/22P docking.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.